# Neural Empirical Interpolation Method for Nonlinear Model Reduction

### v1

@@ 1. Introduction
In the numerical solution of partial differential equations (PDEs) [33], it is often necessary to study the behavior of the physics-based model corresponding to many different parameter configurations. This is the case, for example, in inverse problems, design, and uncertainty quantification, where physical or geometric properties of the problem need to be investigated. Approximating the solutions of the PDE via high-fidelity simulations usually involves solving a system of equations of large dimensionality, and obtaining many-query and real-time simulations becomes prohibitively expensive. Reduced order modeling (ROM) [4, 5] consists of a wide class of approaches in which such solutions are approximated via surrogate models with lower dimensional systems of equations. Because of this reduced or latent representation, the system can be solved more efficiently, allowing for a reduction of the computational cost.
ROM techniques are divided into an offline stage and an online stage, where the information collected via high fidelity solvers during the former is exploited for fast online evaluations. This is possible thanks to the reduced basis space, a low-dimensional space spanned by an orthonormal basis which can be obtained via well-known approaches such as the Greedy method or the Proper Orthogonal Decomposition (POD) [19, 32]. Within the POD-based model order reduction framework, the high fidelity or full-order approximation of the solution is computed and stored offline for several parameter instances, forming the so-called snapshots. Then, a singular value decomposition of the snapshots matrix is performed to retain the hopefully few left singular vectors capturing most of the system’s behavior. A Galerkin projection is then performed on the subspace encoded by the extracted reduced basis, and the high fidelity problem is reformulated in a lower dimensional space. In the online stage, the lower dimensional problem can be solved efficiently and accurately for many different parameter configurations.
To obtain such computational speedup, two main assumptions need to be satisfied: (i) the problem is reducible, meaning that we can compress the information with a small number of basis functions, i.e., its Kolmogorov n-width has a fast decay and (ii) we can efficiently assemble the parameter-dependent quantities thanks to the affine-decomposition property, allowing us to have an online stage that does not depend on the potentially high number of degrees of freedom of the problem. The former is related to the physical complexity of the model, while the latter can be satisfied choosing ad-hoc strategies to restore the computational savings.
@@
For this reason, hyper-reduction techniques have been developed to satisfy the assumption (ii) above, improving the time complexity of evaluating nonlinearities thanks to a further level of model reduction. Among the most common hyper-reduction techniques, there are the Empirical Interpolation Methods (EIM) [3] and its variant the Discrete Empirical Interpolation Methods [8]. In DEIM, one performs an oblique projection onto a subspace of the column space of nonlinearities computed in the offline phase. This projection is chosen in such a way that only a smaller number of entries of the nonlinearity need to be computed in the online phase, and for this reason, it allows for computational cost reduction when dealing with componentwise nonlinearities. For nonlinearities in which the components of the nonlinearity depend on every input to the nonlinearity, exploiting DEIM could not be effective. Moreover, it can only be applied if one has access to the exact form of the nonlinearity in the system.
@@
In this paper, we develop a data-based hyper-reduction machine-learning technique for ROM which we call the Neural Empirical Interpolation Method (NEIM). The idea of NEIM is to approximate an affine decomposition of the nonlinearity in a reduced order model so that the nonlinear term can be written as the sum of the product between parameter-dependent coefficients and solution-dependent neural networks. This approximation is done with a greedy approach. Given a fixed set of parameters, NEIM first finds the parameter corresponding to the reduced order nonlinearity with highest approximate $L^{2}$ error. A neural network is then trained to approximate a normalized version of this nonlinearity, and the coefficient of the first term in the affine decomposition is determined by solving a linear system of equations. This is repeated iteratively, with the $L^{2}$ error computed relative to the current approximation of the affine decomposition. At the end of this iterative procedure, the approximations for the coefficients in the affine decomposition are interpolated from the fixed set of parameters to the whole parameter space. Because of the greedy strategy, we are able to perform a basic error analysis to investigate the performance of NEIM. Namely, we can show that the error estimate used in the algorithm is decreasing with each iteration. We also show how the final generalization error is composed of projection, training, and interpolation errors. NEIM has the advantage over DEIM of being a completely data-driven approach. Moreover, unlike DEIM which involves a linear projection and deals with componentwise nonlinearities, NEIM is a nonlinear methodology which is also suitable for nonlocal nonlinearities.
We demonstrate the effectiveness of our methodology on solution-independent and solution-dependent nonlinearities, a nonlinear elliptic problem, and a nonlinear parabolic model of liquid crystals. In the solution-independent test case, we use NEIM to approximate a nonlinearity which depends only on the PDE parameter $\mu$ and not the PDE solution. In this case, the neural networks in the approximation should be approximately constant. In the solution-dependent test case, we modify the nonlinearity in the first example so that the neural networks should learn non-constant functions. Each of these first two methods are based on a finite difference method. The third example presents an application of NEIM to a nonlinear elliptic problem which is solved with a physics-informed neural network (PINNs) [34] in the ROM context [9]. We then present an application of the methodology to a system of PDEs describing the dynamics of liquid crystal molecules which is discretized with finite elements and has two nonlinearities to approximate.
@@
In addition to the advantages of the proposed methodology, NEIM can easily be used in combination with automatic differentiation in deep learning libraries like PyTorch [29]. This is unlike DEIM in finite element settings in which the assembly of the nonlinearity is not easy to interface between deep learning libraries and finite element libraries. We show the compatibility of NEIM with automatic differentiation in our example with a physics-informed neural network. Furthermore, without relying on a complete approximation of the nonlinear field, NEIM provides more control and interpretability, also thanks to the developed error analysis. Lastly, these architectures are by definition small and fast to train and evaluate, allowing for their usage in standard reduction settings.
@@
There exist several machine learning approaches to model order reduction for parameterized PDEs which tackle the hyper-reduction issue from a different perspective. These can be classified into linear and nonlinear approaches, depending on the maps exploited to approximate the solution field. As an example, neural networks [20, 30, 9], interpolation [7, 12], and Gaussian process regression [18, 21] have been used to map PDE parameters to coefficients of a linear basis for a reduced order subspace. The focus of such works, however, is on approximating the entire parameter-solution mapping corresponding to the reduced order problem and not on designing a hyper-reduction technique to be used online. On the other hand, nonlinear compression has gained much interest to deal with, e.g., advection-dominated problems [13, 38, 39, 22], and more generally to overcome the Kolmogorov barrier [28] by exploiting autoencoder architectures [24, 35, 15, 31]. Since the complex parameter-solution mapping in the nonlinear setting is usually obtained through deep neural networks, the approximated map can be subject to overfitting, require many data to be trained, and lack interpretability. The NEIM approach is also closely related to very recent works exploring the usage of empirical interpolation techniques in combination with neural networks [2, 10], and to the development of adaptive methods of constructing suitable basis functions [1, 14, 16, 11].
Lastly, there are more general machine learning approaches to approximating nonlinear operators in scientific computing [23, 6]. In particular, DeepONets are a popular method used for this purpose [26]. In the most basic setting, DeepONets can be thought of as having one neural network which takes in a parameter and another network which takes the evaluations of a function at various points in space. The output of the DeepONet is the dot product of the outputs of these two networks.
From this perspective, NEIM can be thought of structurally as a specific instance of DeepONet. The difference between NEIM approximations and DeepONets is the way in which NEIM is fit to data in the offline phase. NEIM relies on a greedy method of selecting parameters corresponding to individual terms in the approximation sum. In effect, it is a greedy method of producing a sequence of progressively more accurate DeepONets.
@@
The remainder of the paper is organized as follows. Section 2 presents the reduced order modeling framework which motivates NEIM. Section 3 defines and discusses the NEIM algorithm, and Section 4 presents numerical results. Finally, Section 5 discusses conclusions.
@@ 2. Problem Setting
We begin by describing our setting of reduced order models. Let $p,d\in\mathbb{N}$, and consider $\mathcal{P}\subset\mathbb{R}^{p}$ a compact parameter space and $\Omega(\mu)\subset\mathbb{R}^{d}$ a bounded spatial domain, where $\mu\in\mathcal{P}$. Then for $\mu\in\mathcal{P}$, consider the nonlinear partial differential equation
@@
where $\mathcal{A}(\mu)$ is a linear differential operator, $\mathcal{N}(\mu)$ is a nonlinear differential operator, and $f(\mu)$ is a forcing term. $\mathcal{B}(\mu)$ is a boundary operator defined on the boundary $\partial\Omega(\mu)$ of the spatial domain. We seek a solution $v\in X$, where $X$ is some suitable function space.
To solve this problem numerically, we introduce a finite dimensional subspace $X_{h}\subseteq X$ in which we seek a solution to a weak formulation of Equation (1). Using, for example, a finite element discretization of the above problem, we obtain the nonlinear system of equations (with a slight abuse of notation)
@@ 2.1. Proper Orthogonal Decomposition
Now we describe how we perform the proper orthogonal decomposition to obtain a reduced order modeling corresponding to the discrete problem in Equation (2).
Suppose that we have some number $\nu$ of parameters $\mu_{1}^{POD},\mu_{2}^{POD},\dots,\mu_{\nu}^{POD}$. Then to each parameter $\mu_{i}^{POD}$, there corresponds a solution $v(\mu_{i}^{POD})$ of (2). We collect these solutions into the snapshot matrix
$S=\begin{bmatrix}\vline&&\vline\\ \\[-10.0pt] v(\mu_{1}^{POD})&\dots&v(\mu_{\nu}^{POD})\\ \\[-10.0pt] \vline&&\vline\end{bmatrix}.$
The proper orthogonal decomposition then consists in taking the singular value decomposition of $S=U\Sigma V^{T}$, where $U\in\mathbb{R}^{n\times n}$, $\Sigma\in\mathbb{R}^{n\times\nu}$, and $V\in\mathbb{R}^{\nu\times\nu}$. We take $U_{r}$ to be the matrix consisting of the first $r$ columns of $U$. Then an approximate solution of Equation (2) is found by solving the reduced order problem
@@
so that the solution $v(\mu)$ to Equation (2) and the solution $\tilde{v}(\mu)$ to Equation (3) satisfy $v(\mu)\approx U_{r}\tilde{v}(\mu)$.
Assuming that $A(\mu)$ and $f(\mu)$ admit affine decompositions so that we can write
$A(\mu)=\sum_{i=1}^{p}\theta_{i}^{A}(\mu)A_{i},\quad f(\mu)=\sum_{i=1}^{p}\theta_{i}^{f}(\mu)f_{i}$
for some coefficients $\theta_{i}^{A}(\mu)$ and $\theta_{i}^{f}(\mu)$ and constants $A_{i}\in\mathbb{R}^{n\times n}$ and $f_{i}\in\mathbb{R}^{n}$, we see that the time complexity of evaluating $U_{r}^{\top}A(\mu)U_{r}$ and $U_{r}^{\top}f(\mu)$ is independent of $n$ and only dependent on $r$ since $U_{r}^{\top}A_{i}U_{r}$ and $U_{r}^{\top}f_{i}$ can be pre-computed. However, the time complexity of computing $U_{r}^{\top}\operatorname{N}(U_{r}\tilde{v};\mu)$ still depends on $n$, so we do not see significant improvements in compute time by solving Equation (3) instead of Equation (2). A typical approach to this problem is to use the discrete empirical interpolation method [8]. Here, we instead introduce an iterative method we call the neural empirical interpolation method in which we approximate an affine decomposition of this nonlinear term using a machine learning approach.
@@ 2.2. Discrete Empirical Interpolation Method
We now briefly recall the discrete empirical interpolation method for later comparison. In DEIM, one approximates
@@
where $V_{k}$ is the matrix of the $k$ left singular vectors of the matrix of nonlinear snapshots of the full-order system,
$S_{\text{N}}=\begin{bmatrix}\vline&&\vline\\ \\[-10.0pt] \operatorname{N}(v(\mu_{1}^{POD});\mu_{1}^{POD})&\dots&\operatorname{N}(v(\mu_{\nu}^{POD});\mu_{\nu}^{POD})\\ \\[-10.0pt] \vline&&\vline\end{bmatrix},$
and $P=\begin{bmatrix}e_{p_{1}}&\dots&e_{p_{k}}\end{bmatrix}$ is a matrix with $e_{p_{i}}=\begin{bmatrix}0&\dots&0&1&0&\dots&0\end{bmatrix}^{\top}\in\mathbb{R}^{n}$ the $p_{i}$-th column of the identity matrix for $i=1,\dots,k$. The indices $p_{1},\dots,p_{k}$ are determined by the greedy Algorithm (1). Following the original DEIM paper [8], the notation max is MATLAB notation so that $[|\rho|,p_{\ell}]=\texttt{max}\{|r|\}$ implies $|\rho|=|r_{p_{\ell}}|=\max_{i=1,\dots,n}\{|r_{i}|\}$ with the smallest index taken in case of a tie.

### v2

@@ 1. Introduction
In the numerical solution of Partial Differential Equations (PDEs) [36], it is often necessary to study the behavior of the physics-based model corresponding to many different parameter configurations. This is the case, for example, in inverse problems, design, and uncertainty quantification, where physical or geometric properties of the problem need to be investigated. Approximating the solutions of the PDE via high-fidelity simulations usually involves solving a system of equations of large dimensionality, and obtaining many-query and real-time simulations becomes prohibitively expensive. Reduced order modeling (ROM) [5, 6] consists of a wide class of approaches in which such solutions are approximated via surrogate models with lower dimensional systems of equations. Because of this reduced or latent representation, the system can be solved more efficiently, allowing for a reduction of the computational cost. ROM techniques are divided into an offline stage and an online stage, where the information collected via high fidelity solvers during the former is exploited for fast online evaluations. This is possible thanks to the reduced basis space, a low-dimensional space spanned by an orthonormal basis which can be obtained via well-known approaches such as the Greedy method or the Proper Orthogonal Decomposition (POD) [21, 35]. Within the POD-based model order reduction framework, the high fidelity or full-order approximation of the solution is computed and stored offline for several parameter instances, forming the so-called snapshots. Then, a singular value decomposition of the snapshots matrix is performed to retain the hopefully few left singular vectors capturing most of the system’s behavior. A Galerkin projection is then performed on the subspace encoded by the extracted reduced basis, and the high fidelity problem is reformulated in a lower dimensional space. In the online stage, the lower dimensional problem can be solved efficiently and accurately for many different parameter configurations.
To obtain such computational speedup, two main assumptions need to be satisfied: (i) the problem is reducible, meaning that we can compress the information with a small number of basis functions, i.e., its Kolmogorov n-width has a fast decay, and (ii) we can efficiently assemble the parameter-dependent quantities thanks to the affine-decomposition property, allowing us to have an online stage that does not depend on the potentially high number of degrees of freedom of the problem. The former is related to the physical complexity of the model, while the latter can be satisfied choosing ad-hoc strategies to restore the computational savings.
@@
For this reason, hyper-reduction techniques have been developed to satisfy the assumption (ii) above, improving the time complexity of evaluating nonlinearities thanks to a further level of model reduction. Among the most common hyper-reduction techniques, there are the Empirical Interpolation Methods (EIM) [3] and its variant the Discrete Empirical Interpolation Methods [9]. In DEIM, one performs an oblique projection onto a subspace of the column space of nonlinearities computed in the offline phase. This projection is chosen in such a way that only a smaller number of entries of the nonlinearity need to be computed in the online phase, and for this reason, it allows for computational cost reduction when dealing with componentwise nonlinearities. For nonlocal nonlinearities, in which the components of the nonlinearity depend on every input, e.g. integral kernels or other integral operators like the Boltzmann collision operator [15], exploiting DEIM would not be effective. Moreover, it can only be applied if one has access to the exact form of the nonlinearity in the system, and if this nonlinearity is not programmed to be compatible with automatic differentiation [4], it is not easily used in workflows requiring deep learning. This is the case, for example, in existing reduced order modeling libraries like RBniCS [39].
@@
In this paper, we develop a data-based hyper-reduction machine-learning technique for ROM, compatible with automatic differentiation, which we call the Neural Empirical Interpolation Method (NEIM). The idea of NEIM is to approximate an affine decomposition of the nonlinearity in a reduced order model so that the nonlinear term can be written as the sum of the product between parameter-dependent coefficients and solution-dependent neural networks. This approximation is done with a greedy approach. Given a fixed set of parameters, NEIM first finds the parameter corresponding to the reduced order nonlinearity with highest approximate $L^{2}$ error. A neural network is then trained to approximate a normalized version of this nonlinearity, and the coefficient of the first term in the affine decomposition is determined by solving a linear system of equations. This is repeated iteratively, with the $L^{2}$ error computed relative to the current approximation of the affine decomposition. At the end of this iterative procedure, the approximations for the coefficients in the affine decomposition are interpolated from the fixed set of parameters to the whole parameter space. Because of the greedy strategy, we are able to perform a basic error analysis to investigate the performance of NEIM. Namely, we can show that the error estimate used in the algorithm is decreasing with each iteration. We also show how the final generalization error is composed of projection, training, and interpolation errors. NEIM has the advantage over DEIM of being a completely data-driven approach, which gives NEIM the main advantage of being a far less intrusive method and more easily integrated with other deep learning techniques. Indeed, in the online phase of solving a reduced order model with NEIM, the nonlinear terms in the original full order model are never evaluated, unlike DEIM which still evaluates the nonlinearity at selected rows. Thus, the full order model code does not need to be changed to exploit the NEIM strategy. Moreover, unlike DEIM which involves a linear projection and deals with componentwise nonlinearities, NEIM is a nonlinear methodology which is also suitable for the previously described nonlocal nonlinearities.
@@
We demonstrate the effectiveness of our methodology on solution-independent and solution-dependent nonlinearities, a nonlinear elliptic problem, and a nonlinear parabolic model of liquid crystals. In the solution-independent test case, we use NEIM to approximate a nonlinearity which depends only on the PDE parameter $\mu$ and not the PDE solution. In this case, the neural networks in the approximation should be approximately constant. In the solution-dependent test case, we modify the nonlinearity in the first example so that the neural networks should learn non-constant functions. Each of these first two methods are based on a finite difference method. The third example presents an application of NEIM to a nonlinear elliptic problem which is solved with a physics-informed neural network (PINNs) [37] in the ROM context [10]. We then present an application of the methodology to a system of PDEs describing the dynamics of liquid crystal molecules which is discretized with finite elements and has two nonlinearities to approximate. In addition to the advantages of the proposed methodology, NEIM can easily be used in combination with automatic differentiation in deep learning libraries like PyTorch [32]. This is unlike DEIM in finite element settings in which the assembly of the nonlinearity is not easy to interface between deep learning libraries and finite element libraries. We show the compatibility of NEIM with automatic differentiation in our example with a physics-informed neural network. Furthermore, without relying on a complete approximation of the nonlinear field, NEIM provides more control and interpretability, also thanks to the developed error analysis. Lastly, these architectures are by definition small and fast to train and evaluate, allowing for their usage in standard reduction settings.
@@
There exist several machine learning approaches to model order reduction for parameterized PDEs which tackle the hyper-reduction issue from a different perspective. These can be classified into linear and nonlinear approaches, depending on the maps exploited to approximate the solution field. As an example, neural networks [22, 33, 10], interpolation [8, 13], and Gaussian process regression [20, 23] have been used to map PDE parameters to coefficients of a linear basis for a reduced order subspace. The focus of such works, however, is on approximating the entire parameter-solution mapping corresponding to the reduced order problem and not on designing a hyper-reduction technique to be used online. On the other hand, nonlinear compression has gained much interest to deal with, e.g., advection-dominated problems [14, 41, 42, 24], and more generally to overcome the Kolmogorov barrier [31] by exploiting autoencoder architectures [26, 38, 17, 34, 30]. Since the complex parameter-solution mapping in the nonlinear setting is usually obtained through deep neural networks, the approximated map can be subject to overfitting, require many data to be trained, and lack interpretability. The NEIM approach is also closely related to very recent works exploring the usage of empirical interpolation techniques in combination with neural networks [2, 11], and to the development of adaptive methods of constructing suitable basis functions [1, 16, 18, 12]. Lastly, there are more general machine learning approaches to approximating nonlinear operators in scientific computing [25, 7]. In particular, DeepONets are a popular method used for this purpose [28]. In the most basic setting, DeepONets can be thought of as having one neural network which takes in a parameter and another network which takes the evaluations of a function at various points in space. The output of the DeepONet is the dot product of the outputs of these two networks. From this perspective, NEIM can be thought of structurally as a specific instance of DeepONet. The difference between NEIM approximations and DeepONets is the way in which NEIM is fit to data in the offline phase. NEIM relies on a greedy method of selecting parameters corresponding to individual terms in the approximation sum. In effect, it is a greedy method of producing a sequence of progressively more accurate DeepONets.
@@
The remainder of the paper is organized as follows. Section 2 presents the ROM framework which motivates NEIM. Section 3 defines the NEIM algorithm, and Section 4 presents numerical results. Finally, Section 5 discusses conclusions.
@@ 2. Problem Setting
We begin by describing our setting of reduced order models. Let $p,d\in\mathbb{N}$, and consider $\mathcal{P}\subset\mathbb{R}^{p}$ a compact parameter space and $\Omega(\mu)\subset\mathbb{R}^{d}$ a bounded spatial domain, where $\mu\in\mathcal{P}$. Then for $\mu\in\mathcal{P}$, consider the nonlinear PDE
@@
where $\mathcal{A}(\mu)$ is a linear differential operator, $\mathcal{N}(\mu)$ is a nonlinear differential operator, and $f(\mu)$ is a forcing term. $\mathcal{B}(\mu)$ is a boundary operator defined on the boundary $\partial\Omega(\mu)$ of the spatial domain. We seek a solution $v\in X$, where $X$ is some suitable function space. To solve this problem numerically, we introduce a finite dimensional subspace $X_{h}\subseteq X$ in which we seek a solution to a weak formulation of Equation (1). Using, for example, a finite element discretization of the above problem, we obtain the nonlinear system of equations (with a slight abuse of notation)
@@ 2.1. Proper Orthogonal Decomposition
Now we describe how we perform the proper orthogonal decomposition (POD) to obtain a reduced order modeling corresponding to the discrete problem in Equation (2). Suppose that we have some number $\nu$ of parameters $\mu_{1}^{POD},\mu_{2}^{POD},\dots,\mu_{\nu}^{POD}$. To each parameter $\mu_{i}^{POD}$, there corresponds a solution $v(\mu_{i}^{POD})$ of (2). We collect these solutions into the snapshot matrix $S=\begin{bmatrix}v(\mu_{1}^{POD})&\dots&v(\mu_{\nu}^{POD})\end{bmatrix}$. The POD then consists in taking the singular value decomposition of $S=U\Sigma V^{T}$, where $U\in\mathbb{R}^{n\times n}$, $\Sigma\in\mathbb{R}^{n\times\nu}$, and $V\in\mathbb{R}^{\nu\times\nu}$. We take $U_{r}$ to be the matrix consisting of the first $r$ columns of $U$, and approximate the solution of Equation (2) by solving the ROM problem
@@
so that the solution $v(\mu)$ to Equation (2) and the solution $\tilde{v}(\mu)$ to Equation (3) satisfy $v(\mu)\approx U_{r}\tilde{v}(\mu)$. Assuming that $A(\mu)$ and $f(\mu)$ admit affine decompositions such that
$A(\mu)=\sum_{i=1}^{{q_{1}}}\theta_{i}^{A}(\mu)A_{i},\quad f(\mu)=\sum_{i=1}^{{q_{2}}}\theta_{i}^{f}(\mu)f_{i}$
for some coefficients $\theta_{i}^{A}(\mu)$ and $\theta_{i}^{f}(\mu)$, constants $A_{i}\in\mathbb{R}^{n\times n}$ and $f_{i}\in\mathbb{R}^{n}$, and $q_{1},q_{2}\in\mathbb{N}$, we see that the time complexity of evaluating $U_{r}^{\top}A(\mu)U_{r}$ and $U_{r}^{\top}f(\mu)$ is independent of $n$ and only dependent on $r$ since $U_{r}^{\top}A_{i}U_{r}$ and $U_{r}^{\top}f_{i}$ can be pre-computed. However, the time complexity of computing $U_{r}^{\top}\operatorname{N}(U_{r}\tilde{v};\mu)$ still depends on $n$, so we do not see significant improvements in compute time by solving Equation (3) instead of Equation (2). A typical approach to this problem is to use the discrete empirical interpolation method [9]. Here, we instead introduce an iterative method we call the neural empirical interpolation method in which we approximate an affine decomposition of this nonlinear term using a machine learning approach.
@@ 2.2. Discrete Empirical Interpolation Method
We now briefly recall the discrete empirical interpolation method (DEIM), where one approximates
@@
where $V_{k}$ is the matrix of the $k$ left singular vectors of the matrix of nonlinear snapshots of the full-order system $S_{\text{N}}=\begin{bmatrix}\operatorname{N}(v(\mu_{1}^{POD});\mu_{1}^{POD})&\dots&\operatorname{N}(v(\mu_{\nu}^{POD});\mu_{\nu}^{POD})\end{bmatrix}$ and $P=\begin{bmatrix}e_{p_{1}}&\dots&e_{p_{k}}\end{bmatrix}$ is a matrix with $e_{p_{i}}=\begin{bmatrix}0&\dots&0&1&0&\dots&0\end{bmatrix}^{\top}\in\mathbb{R}^{n}$ the $p_{i}$-th column of the identity matrix for $i=1,\dots,k$. The indices $p_{1},\dots,p_{k}$ are determined by the greedy Algorithm (1). Following the original DEIM paper [9], the notation max is MATLAB notation so that $[|\rho|,p_{\ell}]=\texttt{max}\{|r|\}$ implies $|\rho|=|r_{p_{\ell}}|=\max_{i=1,\dots,n}\{|r_{i}|\}$ with the smallest index taken in case of a tie.